OpenStudy (anonymous):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y

OpenStudy (kewlgeek555):

Let me see if I can dot his. :3

OpenStudy (anonymous):

The first thing you want to do is pick an equation that you can isolate a variable in. The first equation you can isolate y. y - 3x = 7 ---> y = 3x + 7 now sub 3x + 7 in for y in the 2nd equation 21 + 9x = 3y 21 + 9x = 3(3x + 7) -- distribute 21 + 9x = 9x + 21 -- subtract 9x 21 + 9x - 9x = 21 -- subtract 21 9x - 9x = 21 - 21 0 = 0 This system of equations has INFINITE SOLUTIONS

OpenStudy (anonymous):

y - 3x =7 y= 3x + 7. use this statement to substitute. 21 + 9x = 3y (now substitute the previous for y) 21 + 9x = 3(3x + 7) use distributive method 21 + 9x = 9x + 21 Solve for x 0 = 0 Solution: All real numbers (or infinite solutions)

OpenStudy (anonymous):

i still dont get it

OpenStudy (kewlgeek555):

Okay: First, you have to solve one of the equations for either y or x. Luckily, I already this for you:\[y=7+3x\] For the second step, you should substitute the solution of the fist step to the other equation.\[21+9x=3(7+3x)\] For the third step, you solve this equation which @kelliegirl33 already did. :3

OpenStudy (kewlgeek555):

Ooh, let me explain further @zobobozo

OpenStudy (anonymous):

okk

OpenStudy (kewlgeek555):

So, for the third step, you solve the equation. 21 + 9x = 3(3x + 7) This seems hard, but what you have to do is take out those parenthesis by distributing. This means you have to multiply three to 3x and to 7. So the new equation would come out as... 21 + 9x = 9x + 21 Now, you know that you have to subtract 9x so you can have one number in the right side of the equation. 21 + 9x - 9x = 21 Now you must subtract 2. 9x - 9x = 21 - 21 Well, we know that this is equal to zero: 0 = 0 When you solve a system of equations and you get 0=0, it means that there is an infinite number of solutions. This means that in the coordinate plane, there are two lines on top of each other with the same slope and y-intercept.

OpenStudy (anonymous):

it doesn't have to be 0 = 0 to be infinite solutions. All it has to be is equal. It could be 7 = 7 or -4 = -4 and it would have infinite solutions

OpenStudy (kewlgeek555):

Sorry, what she said. ^

OpenStudy (anonymous):

what